OpenStudy (anonymous):

The point (2, 1)is a solution to which of the following equations

OpenStudy (anonymous):

Please help

OpenStudy (eli_moses):

are there any answer choices???

OpenStudy (anonymous):

Yes just wait a minute please;)

OpenStudy (eli_moses):

ok

OpenStudy (anonymous):

A.2x+y=5 and -2+y=2 B.-2-y=-5 and 2x-y=2 C.-x+y=3 and 3x+y=-5 D.x+y=3 and 3x-y=5

OpenStudy (anonymous):

Hello @Eli_Moses are you still there

OpenStudy (noseboy908):

If @Eli_Moses is unable to help for some reason, just mention me and I can.

OpenStudy (anonymous):

ok please help me

OpenStudy (noseboy908):

Ok, sure. To find if a given ordered pair is a solution to a given equation you have to substitute in the x and y values given in the ordered pair (Remember, an ordered pair is "(x,y)") into the given equation(s). So, what x and what y value is given in the original ordered pair?

OpenStudy (anonymous):

The easiest to do (if this was a test) was to plug in the point 2 for x and see if you get -1 as an answer

OpenStudy (anonymous):

1* as an answer

OpenStudy (anonymous):

ok idk what x and y is

OpenStudy (noseboy908):

Ok, just keep in mind what I said about ordered pairs: An ordered pair is, in it's most basic form, "(x,y)". Where x is the x value, and y is the y value. So, what's the ordered pair given in the instructions?

OpenStudy (anonymous):

Ok I think in the first equation the y is 0

OpenStudy (noseboy908):

No, look at the very first thing you said in the instructions. What is the ordered pair given? (You may be right, but i'm not sure where you're getting that y is 0)

OpenStudy (anonymous):

Oh I think the oreed pair I'd 2, q

OpenStudy (anonymous):

Ordered pair is 2,1

OpenStudy (noseboy908):

Correct. So, given what I've told you about how an ordered pair is structured, what is the x value given in it?

OpenStudy (anonymous):

2

OpenStudy (noseboy908):

Yep, now what is the y value?

OpenStudy (anonymous):

1

OpenStudy (noseboy908):

Correct. Now, let's look at the first possible answer. Please restate possible answer A for me.

OpenStudy (anonymous):

2x+y=5 and -2+y=2

OpenStudy (noseboy908):

Yep. So now, let's take the x and y values we get from the given ordered pairs, and plug them in for the x and y in "2x+y=5". Please replace the variables with what the ordered pair's values as I just described, leaving parenthesis around the numbers in place of x and y.

OpenStudy (anonymous):

I don't know what you mean by that

OpenStudy (noseboy908):

Yeah, I went too into detail there admittedly. Allow me to give an example. 

OpenStudy (anonymous):

No the answer would be 7

OpenStudy (noseboy908):

Exactly! Now, that was just to demonstrate what I mean by "plug in the x and y value you're given into the first given equation". Please try to do so.

OpenStudy (anonymous):

Ok 2(2)+1=5 and -2(2)+1=2

OpenStudy (anonymous):

The first equation is true and the second one is not

OpenStudy (noseboy908):

Perhaps you simply mistyped the second equation there, but the way you gave it early on was "-2+y=2". You put a number next to the "-2" in the second equation like there was an equation there.

OpenStudy (anonymous):

Well how am I suppose to do it

OpenStudy (noseboy908):

Either way, let's move on, I believe you understand what you're doing. Let's go to B. If you get what you're doing, please plug in the x and y value where needed into the equations answer B gives us, and list what you get as you just did.

OpenStudy (anonymous):

-2-(1)=-5 and 2(2)-(1)=2

OpenStudy (noseboy908):

Ok, now, are each of those equations true? Work them out if you can't solve them in your head.

OpenStudy (anonymous):

No theyre not true

OpenStudy (noseboy908):

Correct! Now, on to C! Repeat the process please. If you can do it without my help, feel free to just tell me if it does or if it doesn't work out to be true.

OpenStudy (anonymous):

No they're not true also

OpenStudy (anonymous):

And I worked on d and that one is true

OpenStudy (noseboy908):

Indeed! That's it then, D is your answer.

OpenStudy (noseboy908):

Are you still confused about anything?

OpenStudy (anonymous):

No so Thank you very much

OpenStudy (noseboy908):

You are most welcome! I have one request of you before we part: If you don't mind, please leave a testimonial for me. It helps! Thanks!

OpenStudy (anonymous):

Idk whats that

OpenStudy (noseboy908):

Ok, no problem. Start by hovering over either my name or image (left of my replies), and click on "view detailed profile". From there you can leave a testimonial, which tells others how well or bad you think I did as a tutor. Others, if they ever question if I'm worth allowing to tutoring them can look at that and see how you felt I did.

OpenStudy (noseboy908):

All I ask about it is that you're honest. Again, you don't have to if you don't want to, I just appreciate it if you have time to.

OpenStudy (anonymous):

Ok I will try to find and give you a good review

OpenStudy (noseboy908):

Thanks, have a nice evening.

OpenStudy (anonymous):

You to